((( SOLVE 3X^2 -2X + 4 = 0))))) Please help me solve this problem
Answer by mukhopadhyay(490)

3x^2-2x+4=0
Per quadratic formula, roots of ax^2+bx+c are
x=[-b+sqrt(b^2-4ac)]/2a and x=[-b-sqrt(b^2-4ac)]/2a
In this example, a=3, b=-2, and c=4
So, x= [2+sqrt(4-48)]/6 and x= [2-sqrt(4-48)]/6
=> x = [2+sqrt(-44)]/6 and x = [2-sqrt(-44)]/6
=> x = (1+i sqrt(11)/3 and x= (1-isqrt(11)/3